OpenStudy (anonymous):

Find each percent of change and round to the nearest tenth 95 to 76

OpenStudy (anonymous):

30

Directrix (directrix):

19/95 = .2 = 20% decrease

OpenStudy (anonymous):

^ is ok

OpenStudy (anonymous):

yea

Directrix (directrix):

Finding the percent of change is using the ratio of the amount of change to the original amount. The increased amount is really the percent of increase. If the amount decreases then the percent of the change is the percent of the decrease which will be a negative. After you learn to do this type problem, check out the online percentage change calculator. Percentage Change Calculator http://www.csgnetwork.com/percentchangecalc.html

OpenStudy (anonymous):

by comparing the two values, you can get the difference is 76-95= -19 % change= (new value - old value)/old value = -19/95 = -0.2=-20% -ve means the value decreases